# MMDVS-LF: A Multi-Modal Dynamic-Vision-Sensor Line Following Dataset

### v1

@@ III-A Recording
In addition to recording their driving, we asked participants to complete a consent form and a demographic questionnaire. This questionnaire collected their age, gender, country of origin and residence, and health details, including any chronic illnesses, visual impairments, or conditions affecting their vision. We also gathered information about their driving experience, including their length and frequency, professional or racing experience, prior experience with driving F1Tenth cars, comfort level with new technology, and whether they experience motion sickness while driving. Anonymized participants’ data, including the mapping of the recordings to a driver, is available in the raw data.
Figure 2 displays the distribution of selected demographic data. For the Line Following task, we had seven participants, of whom six were born and obtained their driver’s licenses in a country in Western Europe and one in North America. We had six male participants and one female participant, with an age distribution peaking at 25-30, including participants up to 35-40. One participant reported having no or less than one year of experience. Another participant reported 3-5 years of experience; two reported 6-10 years, and three more than ten years. Three participants reported that they have a visual impairment or use a visual aid for driving. Only one participant reported having a chronic illness, which impairs their driving skills, and one participant reported being a professional driver.
TABLE II: Arrays present in a single frame file with their dimensions and a description of their contents. SIZE={512, 256} refers to the resolution size of the dataset.
Name Dimension Description
data (SIZE/2, SIZE, 2) Time until last event since start of frame per channel
mask (SIZE/2, SIZE, 2) Mask for valid data
action (3) Commanded action (steering angle, speed, acceleration)
observation (20) IMU and odometry sensor readings
filtered_mask (SIZE/2, SIZE, 2) Neighborhood filtered mask
owp_mask (SIZE/2, SIZE, 2) Overwrite (clear mask) of earlier events with opposing polarity
filtered_owp_mask (SIZE/2, SIZE, 2) Neighborhood filtered owp_mask
III-B Annotation
We manually annotated the raw data to obtain sections of the recordings with desired behavior. All sections where, the line on the floor is visible in the bottom row of pixels, and where the driver managed to stay on or return to the line without losing it, were considered desired behavior. This extended acceptance leads to a broader range of recorded situations, which should also allow learning-based algorithms to learn recovering behaviors.
We also labeled possible sections that contain objects that have no direct influence on the Line Following task, but might interfere with computer vision applications. Examples are insects detected by the DVS only, and humans standing on or close to the line. The latter occurred in some recordings at the end to mark the end of the recording.
(a) Time surface.
(b) RGB image.
Figure 4: Frame of the data used in the dataset. The time surface is on the left, and the RGB image is on the right. In the time surface, lighter areas indicate earlier events and darker areas later ones. The DVS and RGB camera optics have slightly different intrinsics, leading to different perspectives of the two images.
TABLE III: Number of frames and size of the compressed dataset for the different generated datasets. Values are indicated by the available data with noise / without noise. The 60 Hz version includes RGB data, while the 100 Hz and 120 Hz versions have DVS data only.
Noisy/No Noise 128x256 256x512
60 Hz 136,484 / 96,161 Frames
4.92 GB / 2.926 GB 10.98 GB / 7.462 GB
100 Hz 227,375 / 160,127 Frames
1.56 GB / 1.08 GB 4.98 GB / 3.45 GB
120 Hz 272,838 / 192,127 Frames
1.73 GB / 1.19 GB 5.42 GB / 3.75 GB
Due to differing lighting conditions in the recording area, the infrared lasers of the Intel realsense’s active depth estimation system were visible in some sections of the recordings. They introduced noise into the dataset recording, so we removed the areas with heavy noise from the generated DVS dataset. Some sections contained less noise, which we included in a separate dataset for more robust training.
We derive the action annotations from the human drivers’ driving commands and include observations from some sensors. Other sensors, such as LIDAR, were omitted from the dataset as they are irrelevant to the Line Following task.
III-C Format
From the raw data recorded in Sec. III-A and the annotations, we generated frame-based datasets with frequencies of 60 Hz, 100 Hz, and 120 Hz and image resolutions of 128x256 and 256x512. The dataset with 60 Hz includes RGB images, as we use a camera with 60 FPS for recording. We omitted the RGB images for datasets with higher frequencies to avoid using poor interpolation results. We treat events’ polarity separately for this dataset, generating two channels, one for each polarity.
All *mask arrays represent event frames of the dataset. The data array might contain unfiltered arbitrary data, which must be combined with one *mask array. The action consists of the steering angle and either speed or acceleration commands. The observation array provides data from the IMU sensor, including acceleration in the (x,y,z) directions, angular velocity around these axes, and the orientation quaternion for (x,y,z,w) components. In addition to this, the observation also includes odometry information, such as pose estimation (x,y,z), orientation quaternion for (x,y,z,w), and velocity values along the (x,y,z) axes.
III-D Statistics
@@
We generate twelve datasets with time surfaces and event frames, actions, and observations, based on the different resolutions, frequencies, and the inclusion or exclusion of sections with a small amount of noise. While the representations differ in resolution and generation frequency, the underlying data is the same, and the resulting datasets have the same action distributions. The analysis in this section was performed on the 256x512@60Hz dataset, and light noise sections were included. Other datasets, especially the ones without the light noise sections, might differ slightly.
@@
The generated datasets span 38 minutes, including noisy sections, or 27 minutes without those sections. Depending on the frequency, this leads to datasets of 96,161 to 272,838 frames. Table III gives an overview of the size of the generated datasets and the compressed frames size sum.
@@
We recorded the MMDVS-LF’s data over about 8 hours, including instructions for the drivers, training, setup time, and technically required breaks, such as changing batteries. The annotation of the dataset took approximately four weeks, including regular updates of the annotation tools, as we discovered issues with our tooling or errors in annotation. Generating the dataset with our tooling takes approximately 36 hours on an Intel(R) Xeon(R) Gold 6130 machine with 20 CPU cores and 64 GB of RAM.
IV BENCHMARK
The DVS data holds many promising directions for deep learning research. One can investigate which representation of the data fits better with existing machine learning algorithms or develop new ANN architectures that suit the unique characteristics of DVS data.
Figure 5: Summary showing the different inputs for training deep learning models to predict various outputs using the MMDVS-LF dataset.
TABLE IV: Settings of the layers in the convolutional networks. Adapted from [29].
Layer Type Settings
Input Input shape: (128, 256, 2)
Conv2D Filters: 24, Kernel size: 5, Stride: 2
Conv2D Filters: 36, Kernel size: 5, Stride: 1
MaxPool2D Pool size: 2, Stride: 2
Conv2D Filters: 48, Kernel size: 3, Stride: 1
MaxPool2D Pool size: 2, Stride: 2
Conv2D Filters: 64, Kernel size: 3, Stride: 1
MaxPool2D Pool size: 2, Stride: 2
Conv2D Filters: 64, Kernel size: 3, Stride: 1
Flatten -
Dense Units: 64
@@ IV BENCHMARK
TABLE V: Training, validation, and test losses of different architectures on the MMDVS-LF dataset. We found that all ANNs can converge to comparably the same training and validation loss values in the Line Following task, but only CNNs in combination with strong recurrent parts, such as LSTMs and LTCs, can generalize well on unseen data, which can be observed in the Test loss. Results are averaged over three runs.
@@
CNN $17.137\pm 0.572$ $72.567\pm 4.076$ $197.941\pm 16.456$
CNN + Simple RNN $16.090\pm 0.499$ $75.223\pm 0.437$ $197.605\pm 12.809$
CNN + MGU $15.790\pm 0.740$ $69.077\pm 2.009$ $196.521\pm 20.303$
CNN + GRU $15.187\pm 0.745$ $74.633\pm 1.095$ $180.543\pm 12.349$
CNN + LSTM $22.970\pm 7.953$ $66.963\pm 3.516$ $106.961\pm 6.078$
CNN + LTC $18.99\pm 6.015$ $74.923\pm 4.344$ $122.298\pm 8.078$
IV-A Setups from the dataset
As summarized in Sec. II-B, most machine learning work on DVS data focused on optical flow and object recognition tasks. Here, we aim to highlight various other possibilities our MMDVS-LF dataset provides, such as control tasks (regression), driver identification (classification), and other data science tasks. Fig. 5 summarizes possible training setups using the dataset, including regression tasks predicting the steering angle with velocity or acceleration values based on various combinations of inputs. For classification tasks, one can consider the different drivers completing the Line Following task as class labels and use the available input data (excluding the demographic information) to make the prediction. Suppose someone aims to pursue a data science project. In that case, exploring the correlation between driving characteristics and demographic information or fault detection from the various sensor readings is possible.
IV-B Steering prediction from time surfaces
Here, we present a use case for the MMDVS-LF dataset of 128x256@100Hz, where the goal is to train neural network models to predict the steering angle based on the time surface data from the DVS sensor. As pointed out in Sec. III-D, most of the velocity values fall into a small range, allowing us to simplify the task by treating the speed as constant. The pipeline is illustrated in Fig. 6. We provide the code of a TensorFlow dataloader pipeline and training scripts.
Figure 6: For benchmarking, we use the following architecture: sequences of time surface data are created and fed into our neural networks of interest. These networks consist of several convolutional and max pooling layers, followed by a dense or recurrent network, which predicts the sequence of steering angles corresponding to the input.
We trained a convolutional-neural-network (CNN) front-end [19], with either a fully-connected dense layer, or a Recurrent Neural Network (RNN), respectively, as the back-end policy. As an RNN we used either a fully-connected simple RNN [30], a Minimal Gated United (MGU) [31], a Gated Recurrent Unit (GRU) [32], a Long-Short Term Memory (LSTM) [33], or a Liquid-Time Constant Network (LTC) [34], respectively. In these architectures, the CNN extracts visual information, while the RNN component leverages the sequential nature of the task. For configuring the CNN layers, we adapted the settings from the convolutional head used in [29], which was designed to explore the task of curvature prediction based on RGB images using a combination of CNNs and bio-inspired recurrent models. This adaptation is appropriate because, at a high level, our task is similar from an ML perspective.
We compute the mean squared error (MSE) between the predicted steering angle and the ground truth values over the sequences, scaling the errors by $10^{4}$ for better readability. The data is split into training, validation and test sets with a ratio of 60%/20%/20%. We did hyperparameter tuning for the learning rate in the range of $\{0.0001,0.001,0.01\}$. Based on the best validation loss, we train all networks using a learning rate of $0.0001$. During the training, we use the AdamW optimizer [35] with a cosine weight decay of $10^{-6}$. We run the training for 50 epochs and save the final models with the best validation loss.
The results of these experiments are shown in Table V. We found that all architectures were able to adapt to the task. Our results demonstrate that more sophisticated architectures incorporating recurrent networks generalized better on the MMDVS-LF dataset, leading to smaller loss values.
@@ IV-B Steering prediction from time surfaces
Here, we presented a simple setup from our MMDVS-LF dataset with a wide range of deep-learning approaches. This setup can be extended by using additional available information, such as stacking RGB channels to DVS as extra input channels to the CNN part, resulting in 5 channels (3 channels from RGB, two channels from DVS) in total, and mapping other sensor information of IMU and odometry to the dense or recurrent part. One can extend the output to making sequential predictions not only on the steering angle but also on the velocity or acceleration commands. In this case, one should adapt the loss function to $L=w_{s}\text{MSE}(y_{s},\hat{y}_{s})+w_{v}\text{MSE}(y_{v},\hat{y}_{v})$ to properly scale the mean squared errors of the used commands between the ground truth labels $y_{s},y_{v}$ and predictions $\hat{y}_{s},\hat{y}_{v}$ by the corresponding weights $w_{s}$,$w_{v}$, for the steering and velocity, respectively.
V CONCLUSIONS
@@
We introduced MMDVS-LF, a multimodal, compact, and easy-to-use dataset primarily intended for basic research, focusing on novel deep learning solutions leveraging sparse DVS data for control applications. The paper described the methods for recording experiments and constructing the dataset. We also showed several use cases of our dataset and demonstrated the power of recurrent neural networks predicting steering commands from time surface data.
@@
In the future, we aim to explore DVS-specific control solutions and verify the attention maps of trained neural networks with the recorded eye-tracking data. The relatively inexpensive standardized platform of F1Tenth cars holds the potential to deploy end-to-end machine learning solutions on hardware, making it accessible to universities, research institutions, and the general public to test their solution developed and trained on the MMDVS-LF dataset.
@@ ACKNOWLEDGMENT
We thank Mihaela-Larisa Clement for helping with the data collection and the participants in our recordings.

### v2

@@ IV-E Statistics
We generate 16 datasets with time surfaces and event frames, actions, and observations based on the different resolutions, frequencies, and the inclusion or exclusion of sections with sunlit areas, optionally including the raw events per time frame. While the representations differ in resolution and generation frequency, the underlying data is the same, and the resulting datasets have the same action distributions. The analysis in this section was performed on the 256x512@100Hz dataset, and sunlit sections were included. Other datasets, especially the ones without the sunlit sections, might differ slightly.
@@
The generated datasets span 1:35:52 hours, including sections with sunlit areas, or 1:23:27 hours without those sections. Depending on the frequency, this leads to datasets up to 575,213 frames for the dataset generated at 100Hz with sunlit areas.
@@
We recorded the MMDVS-LF’s data over about 10 hours, including instructions for the drivers, training, setup time, and technically required breaks, such as changing batteries. The annotation of the dataset took approximately two weeks and generating the dataset with our tooling took approximately one week on an Intel(R) Xeon(R) Gold 6130 machine with 20 CPU cores and 64 GB of RAM.
As already pointed out, organizing and recording (in Section IV-A), aligning the frames (in Section IV-B), and processing the dataset (in Section IV-C and IV-D) are time-consuming and generally non-trivial tasks. Developing the tools needed for synchronization, alignment, and processing has occurred over the last three years. Also, organizing participants and researchers for recording sessions in a lab with a fully set up recording environment required careful planning and coordination.
V BENCHMARK
DVS data offers many promising directions for deep learning research. First, we focus on the time surface representation for training machine learning models. Then, we highlight the differences in attention provided by DVS data compared to traditional RGB. Finally, we outline further potential of our MMDVS-LF dataset.
V-A Steering Prediction from Time Surfaces
Here, we present a use case for the MMDVS-LF dataset of 128x256@100Hz, where the goal is to train neural network models to predict the steering angle based on the time surface data from the DVS sensor. As pointed out in Sec. IV-E, most of the velocity values fall into a narrow range, allowing us to simplify the task by treating the speed as constant. The pipeline is illustrated in Fig. 5. We provide the code of a TensorFlow dataloader pipeline, and training and evaluation scripts in our GitHub repository.
Figure 5: For benchmarking, we use the following architecture: sequences of time surface data are created and fed into our neural networks. These networks consist of a CNN Block with several convolutional and max pooling layers, followed by a flattening layer. These features are fed into a fully-connected RNN, which predicts the sequence of steering commands corresponding to the input. Before and after the RNN block, we use additional input and output mappings. For analysis, we apply the VisualBackProp method to extract the attention maps of the trained models. These are compared to the human attention from the eye-tracking data available in our dataset.
We trained a CNN head [25] with an RNN as a policy. As an RNN we used either a fully-connected simple RNN [26], a Minimal Gated Unit (MGU) [27], a Gated Recurrent Unit (GRU) [28], a Long-Short Term Memory (LSTM) [29], a Liquid-Time Constant (LTC) [30], or a Liquid Capacitance Liquid Resistance (LRC) [31] network, respectively. In these architectures, the CNN extracts visual information, while the RNN component leverages the sequential nature of the task. For configuring the CNN layers, we adapted the settings from the convolutional head used in [32], which was designed to explore the task of curvature prediction based on RGB images using a combination of CNNs and bio-inspired recurrent models. This adaptation is appropriate because, at a high level, our task is similar from an ML perspective.
We compute the mean squared error (MSE) between the predicted steering angle and the ground truth values over the sequences and scale the errors by $10^{4}$ for better readability. The data is split into training, validation and test sets with a ratio of 75%/15%/15%. We did hyperparameter-tuning for the learning rate in the range of $\{0.0001,0.001,0.01\}$. Based on their best validation loss, we train all networks using a learning rate of $0.0001$. During the training phase, we use the AdamW optimizer [33] with a cosine weight decay of $10^{-6}$. We run the training for 30 epochs and save the final models with the best validation loss. They are then tested for 2,500 steps on unseen data.
The results of these experiments are shown in Table III. We found that all architectures were able to adapt to the task except for the CNN with Simple RNN. Our results demonstrate that more sophisticated architectures generalized better on the MMDVS-LF dataset, leading to smaller loss values. This also demonstrates that the proposed CNN head is able to extract the necessary features from the time surface, making it usable for the recurrent controllers. This is further supported by the analysis of the attention in Fig. 6, which are calculated by the VisualBackProp algorithm [34], representing where the networks were focusing during decision-making. Note that all networks reduced the noise from the original input and they differ only in which part of the line ahead they prioritize.
@@ V-A Steering Prediction from Time Surfaces
TABLE III: Training, validation, and test losses of different RNNs using the same CNN head on the MMDVS-LF dataset. We found that CNNs in combination with MGUs, LSTMs and LRCs fit better during training and also generalize well on unseen data, which can be observed in the Test loss. Results are averaged over three seeds.
@@
RNN $411.13\pm 8.10$ $168.13\pm 11.98$ $227.63\pm 29.10$
MGU $130.48\pm 35.58$ $55.92\pm 5.17$ $\mathbf{98.74\pm 7.51}$
GRU $146.42\pm 41.20$ $57.42\pm 8.23$ $111.24\pm 7.57$
LSTM $143.93\pm 33.10$ $64.56\pm 6.45$ $\mathbf{103.60\pm 9.83}$
LTC $278.37\pm 28.07$ $100.33\pm 14.04$ $147.69\pm 35.77$
LRC $64.70\pm 4.30$ $43.48\pm 2.78$ $\mathbf{107.29\pm 44.46}$
Figure 6: Attention maps of the networks, computed using the VisualBackProp algorithm, show that the models focus on different parts of the line. When compared to the human attention in Figure 5, which displays the same frame, we found that GRU, LTC, and LRU primarily focus on the same areas as humans.
V-B Impact of the DVS data on Attention Maps
Understanding how different sensor modalities influence deep learning models is crucial for improving performance in real-world applications. To address this, we investigate the impact of input representation on model attention by comparing RGB and DVS time surfaces. In this study, we trained a CNN+LSTM model separately on each data type and analyzed the resulting attention maps, which is presented in Fig. 7. The results show that with RGB data, the model primarily focuses on the beginning of the bold line while also slightly attending to some surrounding areas. In contrast, when using DVS data, the model eliminates the scattered area in the bottom left of the time surface representation and focuses precisely on the line boundaries, which is the most critical feature for predicting steering. This suggests that DVS input provides a highly informative minimal representation for this task, which the model can utilize effectively.
Furthermore, when comparing these findings with the different attention maps in Fig. 6, we observe that the choice of the model significantly influences attention. Some models focus more effectively on the same parts of the curve as our human participants.
To the best of our knowledge, this is the first work to analyze attention maps of DVS data, providing new insights into how event-based vision influences model interpretability and decision-making.
Figure 7: Demonstration of the effectiveness of DVS time surface data. The resulting attention maps show that the network focuses very precisely on the line boundaries when using DVS data, highlighting the advantage of its input representation.
V-C Other Setups from the MMDVS Dataset
@@
In this section, we aim to highlight various possibilities our MMDVS-LF dataset provides, such as control tasks (regression), driver identification (classification), and other data science tasks. In Section V-A, we presented a possible setup from our MMDVS-LF dataset with a wide range of deep learning approaches. This setup can be extended by using additional available information, such as stacking RGB channels to DVS as extra input channels to the CNN part, resulting in 5 channels (3 channels from RGB, two channels from DVS) in total, and mapping other sensor information of IMU and odometry to the dense or recurrent part. One can extend the output to making sequential predictions not only on the steering angle but also on the velocity or acceleration commands. In this case, one should adapt the loss function to $L=w_{s}\text{MSE}(y_{s},\hat{y}_{s})+w_{v}\text{MSE}(y_{v},\hat{y}_{v})$ to properly scale the mean squared errors of the used commands between the ground truth labels $y_{s},y_{v}$ and predictions $\hat{y}_{s},\hat{y}_{v}$ by the corresponding weights $w_{s}$,$w_{v}$, for the steering and velocity, respectively.
There are many possible training setups using the dataset, including classification tasks too, one can consider the different drivers completing the Line Following task as class labels and use the available input data (excluding the demographic information) to make the prediction. Suppose someone aims to pursue a data science project. In that case, exploring the correlation between driving characteristics and demographic information or fault detection from the various sensor readings is possible.
By pursuing any of these directions, new ANN models could be developed to maximize the benefits of DVS data sparsity while also allowing for the integration of eye-tracking data into the training or validation pipeline.
VI CONCLUSIONS
@@
We introduced MMDVS-LF, a multimodal, compact, and easy-to-use dataset primarily intended for basic research, focusing on novel deep learning solutions leveraging sparse DVS and eye-tracking data for control applications. The paper described the methods for recording experiments and constructing the dataset. We also showed several use cases of our dataset and demonstrated the power of RNNs predicting steering commands from time surface representation, and validated their attention by the eye-tracking data.
@@
The relatively inexpensive standardized platform of roboracer cars holds the potential to deploy end-to-end machine learning solutions on hardware, making it accessible to universities, research institutions, and the general public to test their solution developed and trained on the MMDVS-LF dataset.
@@ ACKNOWLEDGMENT
We thank Mihaela-Larisa Clement, Andreas Brandstätter and Moritz Christamentl for helping with the data collection and the participants in our recordings.